# Fast constrained sampling in pre-trained diffusion models

### v2

@@ 3.1 Approximate Newton steps
Lastly, the proposed update of Eq. (10) only requires us to provide the direction of the error ${\bm{e}}$, which points locally towards a lower cost for our constraint on ${\bm{x}}_{0}$. As already discussed, for linear constraints such as inpainting, this direction can be obtained in closed form using the inverse of the operator expressed by ${\bm{A}}$. For non-linear constraints, such as style-guidance, ${\bm{e}}$ can be computed using backpropagation through the differentiable non-linear constraint model. Note that this backpropagation operation does not backpropagate through the denoiser itself, but only through the constraint w.r.t. the estimated $\hat{{\bm{x}}}_{0}$. A general version of the proposed algorithm for linear and non-linear constraints is presented in Algorithm 1. We choose DDIM as our sampling algorithm (Song et al., 2020). Applying other sampling methods should be intuitive by interleaving the diffusion and our gradient updates on ${\bm{x}}_{t}$.
Algorithm 1 The proposed algorithm for constrained sampling using a pre-trained diffusion model.

### v3

@@ 3 Method: Approximate Newton steps
Lastly, the proposed update of Eq. (10) only requires us to provide the direction of the error $\bm{e}$}, which points locally towards a lower cost for the constraint on ${\bm{x}}_{0}$. For linear constraints such as inpainting, this direction can be obtained in closed form using the inverse of the operator expressed by $\bm{A}$}. For non-linear constraints, $\bm{e}$} does not have to be computed in closed form from ${\bm{J}}_{f}^{-1}$, and can be approximated. In our experiments, we use backpropagation through the differentiable non-linear VAE decoder and constraint, which is cheaper than backpropagating through the denoiser.
An alternative would be to utilize another inexact Newton approximation to avoid backpropagating through the constraint altogether. We discuss this in Appendix A.4, where for linear constraints applied to the the VAE output space (i.e. pixels), we propose an inexact Newton method that utilizes the Jacobian of the VAE encoder to approximate the inverse of the VAE decoder Jacobian. Similar approximations have been discussed in the literature before [36] and can be particularly useful when backpropagation is infeasible [41].
The proposed method for sampling with linear and non-linear constraints is described in Algorithm 1, using DDIM as the diffusion sampling algorithm [33]. Using other diffusion sampling algorithms [19] is intuitive by interleaving the diffusion and our constraint gradient updates on ${\bm{x}}_{t}$ (Appendix A.6).
Algorithm 1 The proposed algorithm for sampling under linear and non-linear constraints.